OpenStudy (anonymous):

A square has an area of 16 square millimeters. What is the length of each side of the square?

OpenStudy (liljj1421):

4

OpenStudy (anonymous):

The area of a parallelogram is 44.4 square feet and its height is 7.4 feet. What is the length of the base? Do not round your answer.

OpenStudy (liljj1421):

A=b*h 16=b*h 16/4=4 (because of the 4 sides you divide by 4)

OpenStudy (liljj1421):

the formula for area of a parallelogram is A=b*h

OpenStudy (liljj1421):

you already know the area(44.4) and the base (7.4)

OpenStudy (liljj1421):

you would take 44.4 and divide it by the height(7.4) and then you get your answer